1. Given the demand function P= 20-5Q, find the price elasticity of demand when price of the commodity is 5 Birr per unit. Mention if the demand is price elastic or inelastic at this point


1
Expert's answer
2021-08-30T12:34:57-0400

"P= 20-5Q"

"5 = 20 - 5Q"

"5 - 20 = -5Q"

"-15 = -5Q"

"Q = 3"

"Ed=" "dQ\/dP*P\/Q"

"Ed=-1\/5*5\/3 = -0.3333"

Since the value of the elasticity is negative, The demand is price inelastic.